OpenStudy (sophadof):

f(x)=x^2+x-2 and g(x)=2x-4 Identify (f+g)(x)

OpenStudy (mustbethestars):

So it's asking you to add. x^2 + x - 2 + 2x - 4

OpenStudy (mustbethestars):

Combine like terms and add/subtract them tell me what you get.

OpenStudy (michele_laino):

by definition, we can write this: \[\left( {f + g} \right)\left( x \right) \equiv f\left( x \right) + g\left( x \right) = \left( {{x^2} + x - 2} \right) + \left( {2x - 4} \right) = ...?\]

OpenStudy (anonymous):

the answer is f(x+g) = x^2 + 3x -6

OpenStudy (mustbethestars):

Good Job. Welcome to OpenStudy. But as a future reference please do not hand out direct answers.